A cube and a sphere,with equal mass of 4.0kg separated by a 2.0m distance,are shown.If the length of the cube is 1.0 m, and the diameter of the sphere is 1.2m, what is the gravitational force between the two


1
Expert's answer
2021-05-08T14:40:07-0400
"F=G\frac{m^2}{r^2}\\F=6.67\cdot10^{-11}\frac{4^2}{2^2}=2.7\cdot10^{-10}N"